y" — 9y = 0. + e.
1
Expert's answer
2020-12-22T17:19:27-0500

"\frac{d^2y(x)}{dx^2}-9y(x)=e"

The general solution will be the sum of the complimentary solution and particular solution.

The complimentary solution:

"\frac{d^2y(x)}{dx^2}-9y(x)=0"

Solution will be proportional for ekx for some constant k.

Substitute y(x)= ekx into equalation.

k2ekx-9ekx=0

ekx(k2-9)=0

As ekx can not be equal 0, so k2-9=0

k1=3, k2=-3

The complimentary solution is

y(x)=C1e3x+C2e-3x

Particular solution:

y(x)=Be3x

y''=9Be3x

-9Be3x=e

"B=\frac{-e}{9e^{3x}}"

y(x)=-e/9

Answer:y(x)=C1e3x+C2e-3x-e/9